OpenStudy (anonymous):

In the products 2axla =4 1a ; 2bx1b=37b, a and b represent positive integers. Then a+b=

OpenStudy (jack1):

2a x a = 41a ; 2b x b = 37b is that right...?

OpenStudy (jack1):

so 2a^2 = 41a EQN1 and 2b^2 = 37b ENQ2 therefore : 2a = 41 and 2b = 37 so a = 41/2 and b = 37/2 so now add a + b